OpenStudy (anonymous):

Which expression is the simplified form of the square root of 180 m to the fifth

OpenStudy (anonymous):

\[\huge \sqrt{180m^5}=6 \sqrt{5} m^{\frac{5}{2}}\]

OpenStudy (anonymous):

Because: \[\huge \sqrt{x}=x^{\frac {1}{2}}\]

OpenStudy (anonymous):

and 6*6*5 = 180